# Hybrid Type-II and Type-III seesaw model for the muon $g-2$ anomaly

### v1

@@ II.3 Independent input parameters
  • We require $|s^{h}|\lesssim\mathcal{O}(0.1)$ to safely pass the current constraints from the study of Higgs data [49, 50]. In practical numerical calculations we utilize the rich parameter space in $V(H,S)$ to simply set
    $s^{h}=0\,.$ (35)
@@
$\{\,m_{\nu}\,,\quad\,m_{\ell}\,,\,\quad m_{F^{0}}\,\}\,.$ (38)
Similar to Eq. (33), hereafter we would use $m_{F}$ as the notation
@@
to denote the numerical approximations while keeping in mind they can be independent quantities in principle as discussed above. As discussed in Eq. (40), in this work we would take
@@
As for the choice of the other two parameters, given the almost linear correlation shown in Eq. (26), we should not choose them simultaneously. In this work, we consider the following two different schemes.
@@
    In this scheme, Eqs. (22) and (II.2) suggests that the other parameters can be expressed as
@@
    In this scheme, Eqs. (22) and (II.2) suggests that the other parameters can be expressed as
@@
As we will discuss later, the $\{\lambda_{L},x_{L}\}$ and $\{\lambda_{L},\theta_{L}^{\nu}\}$ schemes are physically suitable for the illustration of the decoupling behavior and the illustration of chiral enhancement effects in the $(g-2)_{\mu}$ predictions of our model, respectively.
@@
Fixed $:\,s^{h}=0\,,\quad v_{\delta}=5\,\textrm{GeV}\,,\quad m_{\delta}=1000\,\textrm{GeV}\,,\quad m_{\nu}=0\,,\quad m_{\ell}=105.66\mathrm{MeV}\,,$
Varying $:\,m_{F}\,\qquad\qquad\textrm{ and }\qquad\quad\{\,\lambda_{L},\,x_{L}\,\}\,\textrm{ or }\,\{\,\lambda_{L},\,\theta_{L}^{\nu}\,\}\,.$ (43)
@@ II.4 Perturbativity requirement
  • In the $\{\lambda_{L},\theta_{L}^{\nu}\}$ scheme, we have the following approximations from Eq. (II.3):
@@ III.1 Analytical results
In Fig. 1 we show the Feynman diagrams contributing to the $(g-2)_{\mu}$ in our model. They are divided into three classes according to the fermions appearing in the loop. Originating from the left, middle and right panel, respectively, the analytical results of $a^{\textrm{NP}}$ generated by our model can be decomposed to
$a^{\textrm{NP}}_{\mu}\equiv a_{\mu}^{\ell,\,\textrm{total}}+a_{\mu}^{\nu,\,\textrm{total}}+a_{\mu}^{F,\,\textrm{total}}\,.$ (47)
Note that in Fig. 1 the leptons $\ell,\nu$ should be understood to carry specifically the muon flavor as $\mu,\nu_{\mu}$
Figure 1: Feynman diagrams contributing to the $(g-2)_{\mu}$ in which the left, middle and right panel corresponds to $a_{\mu}^{\ell,\,\textrm{total}},\,a_{\mu}^{F,\,\textrm{total}},\,a_{\mu}^{\nu,\,\textrm{total}}$ in Eq. (47), respectively. Note that in left and middle diagrams the photon can be emitted from either of the two charged particles in the loop, while the emission only takes place from $\delta^{-}$ in the right diagram.
To write down our analytical results, we first define the following expressions as the reduced form of loop functions (see [57] for a more complete list).
@@
$a_{\mu}^{\ell,\,\textrm{total}}$ $=a_{\mu}^{\ell,\,h}+a_{\mu}^{\ell,\,\delta^{0}}+a_{\mu}^{\ell,\,\delta^{--}}\,,$
$a_{\mu}^{\ell,\,h}$ $=\quad\frac{m_{\mu}^{2}}{4\pi^{2}m_{h}^{2}}[\big{(}\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}s_{L}^{\ell}c_{R}^{\ell}s^{h}\big{)}^{2}-\frac{m_{\mu}^{2}}{v^{2}}]\cdot[F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})+F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})]\,,$
$a_{\mu}^{\ell,\,\delta^{0}}$ $=\quad\frac{m_{\mu}^{2}}{4\pi^{2}m_{\delta^{0}}^{2}}[\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}^{\ast}s_{L}^{\ell}c_{R}^{\ell}c^{h}]^{2}\cdot[F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{0}}^{2}})+F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{0}}^{2}})]\,,$
$a_{\mu}^{\ell,\,\delta^{--}}$ $=\quad\frac{m_{\mu}^{2}}{2\pi^{2}m_{\delta^{--}}^{2}}[|x_{L}|^{2}(c_{L}^{\ell})^{4}+|\lambda_{L}|^{2}(s_{R}^{\ell})^{2}(c_{R}^{\ell})^{2}]\cdot[-F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})+2F_{LL}^{S,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})]$
$\quad+\frac{m_{\mu}^{2}}{2\pi^{2}m_{\delta^{--}}^{2}}[-x_{L}\lambda_{L}^{\ast}-\lambda_{L}x_{L}^{\ast}](c_{L}^{\ell})^{2}s_{R}^{\ell}c_{R}^{\ell}\cdot[-F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})]\,.$ (51)
Note that we have subtracted the SM Higgs contribution from $a_{\mu}^{\ell,\,h}$ to meet the definition of $a^{\textrm{NP}}$ generated from new physics sector. Moreover, the symmetry factor from the coupling of ¯$\bar{\ell}\ell^{C}\delta^{--}$ has been properly considered in $a_{\mu}^{\ell,\,\delta^{--}}$ as pointed out in Ref. [58].
@@
$a_{\mu}^{F,\,\textrm{total}}$ $=a_{\mu}^{F,\,h}+a_{\mu}^{F,\,\delta^{0}}+a_{\mu}^{F,\,\delta^{--}}+a_{\mu}^{F,\,\delta^{-}}\,,$
$a_{\mu}^{F,\,h}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{h}^{2}}[(\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}s^{h})^{2}+(\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}-\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}s^{h})^{2}]\cdot F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{h}^{2}}(\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}s^{h})(\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}-\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}s^{h})\cdot F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})\,,$
$a_{\mu}^{F,\,\delta^{0}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{0}}^{2}}[(-\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}c^{h})^{2}+(\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}+\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}c^{h})^{2}]\cdot F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{0}}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{0}}^{2}}(\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}c^{h})(\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}+\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}c^{h})\cdot F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{0}}^{2}})\,,$
$a_{\mu}^{F,\,\delta^{--}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{--}}^{2}}[4|x_{L}|^{2}(s_{L}^{\ell})^{2}(c_{L}^{\ell})^{2}+|\lambda_{L}|^{2}\big{(}(s_{R}^{\ell})^{2}-(c_{R}^{\ell})^{2}\big{)}^{2}]\cdot[-F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LL}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}[x_{L}\lambda_{L}^{\ast}+\lambda_{L}x_{L}^{\ast}]s_{L}^{\ell}c_{L}^{\ell}\big{(}(c_{R}^{\ell})^{2}-(s_{R}^{\ell})^{2}\big{)}\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]\,,$
$a_{\mu}^{F,\,\delta^{-}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{-}}^{2}}[2|x_{L}|^{2}(s_{L}^{\nu})^{2}(c_{L}^{\ell})^{2}+|\lambda_{L}|^{2}(c_{L}^{\nu})^{2}(c_{R}^{\ell})^{2}]\cdot F_{LL}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})$
$\quad+\frac{\sqrt{2}m_{\mu}m_{F^{0}}}{8\pi^{2}m_{\delta^{-}}^{2}}[x_{L}\lambda_{L}^{\ast}+\lambda_{L}x_{L}^{\ast}]s_{L}^{\nu}c_{L}^{\ell}c_{L}^{\nu}c_{R}^{\ell}\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,.$ (52)
@@
$a_{\mu}^{\nu,\,\textrm{total}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{-}}^{2}}[2|x_{L}|^{2}(c_{L}^{\nu})^{2}(c_{L}^{\ell})^{2}+|\lambda_{L}|^{2}(s_{L}^{\nu})^{2}(c_{R}^{\ell})^{2}]\cdot F_{LL}^{S,3}(\frac{m_{\mu}^{2}}{m_{\delta^{-}}^{2}})$
$\quad+\frac{m_{\mu}m_{\nu}}{8\pi^{2}m_{\delta^{-}}^{2}}(-2\sqrt{2}x_{L}\lambda_{L}s_{L}^{\nu}c_{L}^{\nu}c_{L}^{\ell}c_{R}^{\ell})\cdot F_{LR}^{S,3}(\frac{m_{\mu}^{2}}{m_{\delta^{-}}^{2}})\,.$ (53)
One can clearly see that for each one of the four pieces consisting $a_{\mu}^{F,\,\textrm{total}}$, the chiral enhancement indicated by $m_{F^{0}}$ or $m_{F^{\pm}}$ appears in the numerator of the second line. As a result of this enhancement, $a_{\mu}^{F,\,\textrm{total}}$ in Eq. (47) turns out to play the absolutely dominant role compared to $a_{\mu}^{\ell,\,\textrm{total}}$ and $a_{\mu}^{\nu,\,\textrm{total}}$, i.e.
@@
Furthermore, for the different contributions to $a_{\mu}^{F,\,\textrm{total}}$ in Eq. (III.1) , we find that $a_{\mu}^{F,\,\delta^{--}}$ and $a_{\mu}^{F,\,\delta^{-}}$ contribute the dominant and sub-dominant part, respectively.
To simplify our analytical results and highlight the chiral enhancement, we utilize the small mixing condition $s^{\nu}_{L},s^{\ell}_{L},s^{\ell}_{R},s^{h}\ll 1$ discussed in Section II.2 and II.3 and extract the most relevant terms in Eq. (III.1) as follows

### v2

@@ II.3 Input parameters
  • As for the mixing $s^{h},s^{a},s^{G}$ in the scalar sector, we first have the following requirement to safely pass the current constraints from the study of Higgs data [51, 52]
    $|s^{h}|\lesssim\mathcal{O}(0.1)\,,$ (37)
    which can be properly realized due to the rich parameter space of $V(H,S)$. Moreover, our choice of $v_{\delta},v_{h}$ in Eq. (31) and (33) also results in [36]
    $s^{a}\approx\tan\theta^{a}=\frac{2\,v_{\delta}}{v_{h}}\sim\mathcal{O}(10^{-2})\,,\quad s^{G}\approx\tan\theta^{G}=\frac{1}{\sqrt{2}}\tan\theta^{a}\sim\mathcal{O}(10^{-2})\,,$ (38)
    which make the mixing $s^{h},s^{a},s^{G}$ all in the small value region. While keeping $s^{h},s^{a},s^{G}$ in the analytical results, we will simply impose the following simplifications in numerical calculation which make negligible difference to the results
    $s^{h}=s^{a}=s^{G}=0\,.$ (39)
@@
$m_{\nu}\,,\quad\,m_{\ell}\,,\,\quad m_{F^{0}}\,.$ (42)
According to Eq. (28) and similar to Eq. (35), hereafter we would use $m_{F}$ as the following notation
@@
to denote the numerical approximations, while keeping in mind they can be independent quantities in principle as discussed above. As discussed in Eq. (29), in this work we would take
@@
As for the choice of the other two independent parameters, given the almost linear correlation between $s^{\nu}_{L}$ and $s^{\ell}_{L}$ shown in Eq. (26), we should not choose them simultaneously. In this work, we consider the following two different schemes.
@@
    In this scheme, Eqs. (22) and (II.2) suggest that the other parameters can be expressed as
@@
    Again, note that in our numerical analysis we choose all Yukawa couplings to be real and in our conventions we have $x_{L}<0$ and $y,z_{L},\lambda_{L}>0$.
@@
    In this scheme, Eqs. (22) and (II.2) suggest that the other parameters can be expressed as
@@
As we will discuss later, the $\{\lambda_{L},x_{L}\}$ and $\{\lambda_{L},\theta_{L}^{\nu}\}$ schemes are physically suitable for the illustration of the decoupling behavior and chiral enhancement effects in the predictions of $(g-2)_{\mu}$ in our model, respectively.
@@
[…]
$\quad v_{\delta}=5\,\textrm{GeV}\,,\quad m_{\delta}=1000\,\textrm{GeV}\,,\quad\quad s^{h}=s^{a}=s^{G}=0\,,$
[…]
@@ II.4 Perturbativity requirement
  • In the $\{\lambda_{L},\theta_{L}^{\nu}\}$ scheme, we have the following approximations from Eq. (II.3)
@@ III.1 Analytical results
Figure 1: Feynman diagrams contributing to the $(g-2)_{\mu}$ in which the left, middle and right panel corresponds to $a_{\mu}^{\ell,\,\textrm{total}},\,a_{\mu}^{F,\,\textrm{total}},\,a_{\mu}^{\nu,\,\textrm{total}}$ in Eq. (51), respectively. Note that in the left and middle diagrams the photon can be emitted from either of the two charged particles in the loop, while the emission only takes place from $\delta^{-}$ in the right diagram. The Feynman diagrams are drawn by JaxoDraw [59].
In Fig. 1 we show the Feynman diagrams contributing to $(g-2)_{\mu}$ in our model, in which the leptons $\ell,\nu$ should be understood to carry the muon flavor as $\mu,\nu_{\mu}$. They are divided into three parts according to the fermions appearing in the loop. Originating from the left, middle and right panel, respectively, the analytical results of $a^{\textrm{NP}}$ generated by our model can be decomposed to
$a^{\textrm{NP}}_{\mu}\equiv a_{\mu}^{\ell,\,\textrm{total}}+a_{\mu}^{F,\,\textrm{total}}+a_{\mu}^{\nu,\,\textrm{total}}\,.$ (51)
Note that all Yukawa couplings in our model are chosen to be real, but we will present our analytical results in the context of complex Yukawa couplings for the more general scenario. To write down the analytical results, we first define the following expressions as the reduced form of loop functions (see [60] for a more complete list).
@@
$a_{\mu}^{\ell,\,\textrm{total}}=a_{\mu}^{\ell,\,h}+a_{\mu}^{\ell,\,\delta^{0}}+a_{\mu}^{\ell,\,a^{0}}+a_{\mu}^{\ell,\,\delta^{--}}\,,$ (55)
in which the neutral CP-even scalars $h,\delta^{0}$ contribute as
$a_{\mu}^{\ell,\,h}$ $=\quad\frac{m_{\mu}^{2}}{4\pi^{2}m_{h}^{2}}[|\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}c_{R}^{\ell}s^{h}|^{2}-\frac{m_{\mu}^{2}}{v^{2}}]\cdot F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})$
$\quad+\frac{m_{\mu}^{2}}{4\pi^{2}m_{h}^{2}}\textrm{Re}[(\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}c_{R}^{\ell}s^{h})^{2}-\frac{m_{\mu}^{2}}{v^{2}}]\cdot F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{h}^{2}})\,,$
$a_{\mu}^{\ell,\,\delta^{0}}$ $=\quad\frac{m_{\mu}^{2}}{4\pi^{2}m_{\delta^{0}}^{2}}|\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}c_{R}^{\ell}c^{h}|^{2}\cdot F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{0}}^{2}})$
$\quad+\frac{m_{\mu}^{2}}{4\pi^{2}m_{\delta^{0}}^{2}}\textrm{Re}[(\frac{m_{\mu}}{v_{h}}(c_{L}^{\ell})^{2}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}c_{R}^{\ell}c^{h})^{2}]\cdot F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{0}}^{2}})\,,$ (56)
and the neutral CP-odd scalar $a^{0}$ contributes as
$a_{\mu}^{\ell,\,a^{0}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{a^{0}}^{2}}|\lambda_{L}|^{2}(c^{a}s^{\ell}_{L}c^{\ell}_{R})^{2}\cdot F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{a^{0}}^{2}})$
$\quad+\frac{m_{\mu}^{2}}{8\pi^{2}m_{a^{0}}^{2}}\textrm{Re}[-(\lambda_{L}c^{a}s^{\ell}_{L}c^{\ell}_{R})^{2}]\cdot F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{a^{0}}^{2}})\,,$ (57)
and the charged scalar $\delta^{--}$ contributes as
$a_{\mu}^{\ell,\,\delta^{--}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{--}}^{2}}[|x_{L}|^{2}(2c_{L}^{\ell})^{2}+|\lambda_{L}|^{2}(2s_{R}^{\ell}c_{R}^{\ell})^{2}]\cdot[-F_{LL}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})+2F_{LL}^{S,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})]$
$\quad+\frac{m_{\mu}^{2}}{4\pi^{2}m_{\delta^{--}}^{2}}\textrm{Re}[\big{(}-2x_{L}(c_{L}^{\ell})^{2}\big{)}(2\lambda_{L}^{\ast}s_{R}^{\ell}c_{R}^{\ell})]\cdot[-F_{LR}^{f,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,1}(\frac{m_{\mu}^{2}}{m_{\delta^{--}}^{2}})]\,.$ (58)
Note that we have subtracted the SM Higgs contribution from $a_{\mu}^{\ell,\,h}$ to meet the definition of $a^{\textrm{NP}}$, i.e. the contribution exclusively generated from new physics sector. Moreover, the symmetry factor from the coupling of ¯$\bar{\ell}\ell^{C}\delta^{--}$ has been properly considered in $a_{\mu}^{\ell,\,\delta^{--}}$ as pointed out in Ref. [61].
@@
$a_{\mu}^{F,\,\textrm{total}}=a_{\mu}^{F,\,h}+a_{\mu}^{F,\,\delta^{0}}+a_{\mu}^{F,\,a^{0}}+a_{\mu}^{F,\,\delta^{--}}+a_{\mu}^{F,\,\delta^{-}}\,,$ (59)
in which the neutral CP-even scalars $h,\delta^{0}$ contribute as
$a_{\mu}^{F,\,h}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{h}^{2}}[|\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}s^{h}|^{2}+|\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}-\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}s^{h}|^{2}]\cdot F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{h}^{2}}\textrm{Re}[(\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}+\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}s^{h})(\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}c^{h}-\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}s^{h})]\cdot F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{h}^{2}})\,,$
$a_{\mu}^{F,\,\delta^{0}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{0}}^{2}}[|\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}c^{h}|^{2}+|\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}+\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}c^{h}|^{2}]\cdot F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{0}}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{0}}^{2}}\textrm{Re}[(\frac{m_{F^{\pm}}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}-\frac{1}{\sqrt{2}}\lambda_{L}s_{L}^{\ell}s_{R}^{\ell}c^{h})(\frac{m_{\mu}}{v_{h}}s_{L}^{\ell}c_{L}^{\ell}s^{h}+\frac{1}{\sqrt{2}}\lambda_{L}c_{L}^{\ell}c_{R}^{\ell}c^{h})]\cdot F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{0}}^{2}})\,,$ (60)
and the neutral CP-odd scalar $a^{0}$ contributes as
$a_{\mu}^{F,\,a^{0}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{a^{0}}^{2}}[\frac{1}{2}|\lambda_{L}|^{2}(c^{a}s^{\ell}_{L}s^{\ell}_{R})^{2}+\frac{1}{2}|\lambda_{L}|^{2}(c^{a}c^{\ell}_{L}c^{\ell}_{R})^{2}]\cdot F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{a^{0}}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{a^{0}}^{2}}\textrm{Re}[(\frac{1}{\sqrt{2}}\lambda_{L}c^{a}s^{\ell}_{L}s^{\ell}_{R})(\frac{1}{\sqrt{2}}\lambda_{L}c^{a}c^{\ell}_{L}c^{\ell}_{R})]\cdot F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{a^{0}}^{2}})\,,$ (61)
and the charged scalars $\delta^{-},\delta^{--}$ contribute as
$a_{\mu}^{F,\,\delta^{--}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{--}}^{2}}[4|x_{L}|^{2}(s_{L}^{\ell}c_{L}^{\ell})^{2}+|\lambda_{L}|^{2}\big{(}(c_{R}^{\ell})^{2}-(s_{R}^{\ell})^{2}\big{)}^{2}]\cdot[-F_{LL}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LL}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$\quad+\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}\textrm{Re}[(2x_{L}s_{L}^{\ell}c_{L}^{\ell})\big{(}\lambda_{L}^{\ast}((c_{R}^{\ell})^{2}-(s_{R}^{\ell})^{2})\big{)}]\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]\,,$
$a_{\mu}^{F,\,\delta^{-}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{-}}^{2}}[2|x_{L}|^{2}(s_{L}^{\nu}c_{L}^{\ell}c^{G})^{2}+|\lambda_{L}|^{2}(c_{L}^{\nu}c_{R}^{\ell}c^{G})^{2}]\cdot F_{LL}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})$
$\quad+\frac{m_{\mu}m_{F^{0}}}{4\pi^{2}m_{\delta^{-}}^{2}}\textrm{Re}[(\sqrt{2}x_{L}s_{L}^{\nu}c_{L}^{\ell}c^{G})(\lambda_{L}^{\ast}c_{L}^{\nu}c_{R}^{\ell}c^{G})]\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,,$ (62)
@@
$a_{\mu}^{\nu,\,\textrm{total}}$ $=\quad\frac{m_{\mu}^{2}}{8\pi^{2}m_{\delta^{-}}^{2}}[2|x_{L}|^{2}(c_{L}^{\nu}c_{L}^{\ell}c^{G})^{2}+|\lambda_{L}|^{2}(s_{L}^{\nu}c_{R}^{\ell}c^{G})^{2}]\cdot F_{LL}^{S,3}(\frac{m_{\mu}^{2}}{m_{\delta^{-}}^{2}})$
$\quad+\frac{m_{\mu}m_{\nu}}{4\pi^{2}m_{\delta^{-}}^{2}}\textrm{Re}[(-\sqrt{2}x_{L}c_{L}^{\nu}c_{L}^{\ell}c^{G})(\lambda_{L}^{\ast}s_{L}^{\nu}c_{R}^{\ell}c^{G})]\cdot F_{LR}^{S,3}(\frac{m_{\mu}^{2}}{m_{\delta^{-}}^{2}})\,.$ (63)
One can clearly see that for each one consisting $a_{\mu}^{F,\,\textrm{total}}$ in Eq. (59), the chiral enhancement appears as indicated by $m_{F^{0}}$ or $m_{F^{\pm}}$ in the numerator of the second line. As a result of this enhancement, $a_{\mu}^{F,\,\textrm{total}}$ in Eq. (51) turns out to play the absolutely dominant role compared to $a_{\mu}^{\ell,\,\textrm{total}}$ and $a_{\mu}^{\nu,\,\textrm{total}}$, i.e.
@@
Furthermore, for the different contributions to $a_{\mu}^{F,\,\textrm{total}}$ in Eq. (59) , we find that $a_{\mu}^{F,\,\delta^{--}}$ and $a_{\mu}^{F,\,\delta^{-}}$ contribute the dominant and sub-dominant part, respectively.
To simplify our analytical results and highlight the chiral enhancement, we utilize the small mixing condition $s^{\nu}_{L},s^{\ell}_{L},s^{\ell}_{R},s^{h},s^{a},s^{G}\ll 1$ discussed in Section II.2 and II.3 and extract the most relevant terms in Eq. (59) as follows